OpenStudy (anonymous):

deteremine the equation of each line the slope is 6, and (2, -6) is a point on the line

OpenStudy (anonymous):

using the point slope formula y - y1 = m(x - x1) we have y + 6 = 6(x -2) or y + 6 = 6x - 12 or y = 6x - 18

OpenStudy (anonymous):

\[y+6=6(x-2) \implies y-6x=-18 \]